OpenStudy (anonymous):

Anybody know about addition property of equality i dont get it please help :)

hartnn (hartnn):

do you know what addition property of equality is ?

OpenStudy (anonymous):

yea im just confused and i need someone to make it clear for me please

hartnn (hartnn):

addition property of equality : if x=y then adding 'z' to both sides gives x+z=y+z what did u not get in this ?

hartnn (hartnn):

example : if x-3 = y then adding 3 to both sides will give you : x-3+3 = y+3 ---- > x=y+3

OpenStudy (anonymous):

i have an example solve x -4=20 for x

hartnn (hartnn):

add 4 on both sides, what you get ?

OpenStudy (anonymous):

so it would be x-4+4=20 +4 equals 24 ?! :)

hartnn (hartnn):

absolutely correct! good :) x=24.

OpenStudy (anonymous):

thanks i get it now :)

hartnn (hartnn):

welcome ^_^ glad to help :)

OpenStudy (anonymous):

youre a great teacher

OpenStudy (anonymous):

;)

hartnn (hartnn):

thanks a lot :D

OpenStudy (anonymous):

hey im kinda new and how do you give people medals

hartnn (hartnn):

by clicking on 'Best Response' tab , somewhere near here .....................^

hartnn (hartnn):

got that ^ ?

OpenStudy (anonymous):

yeup ;) and thanks i feel welcomed lol

hartnn (hartnn):

feel free to ask about anything else ....

OpenStudy (anonymous):

okay ;)

OpenStudy (anonymous):

is there any way to change your picture from diaganol to horizontal :)

hartnn (hartnn):

not in OS, you have to rotate your pic in your computer, using paint or other software, and then upload it here....

OpenStudy (anonymous):

thanks duh i knew that

OpenStudy (anonymous):

lol ;)